# ConceptPsy: A Benchmark Suite with Conceptual Comprehensiveness in Psychology

### v3

@@ 4.3 Results
In Table 6, we showcase the performance of popular Chinese LLMs. The Yi-34B-Chat model outperforms, achieving an average score of 0.85, surpassing strong models such as Qwen1.5-72B-Chat and GPT-3.5-Turbo. Both Qwen1.5-14B-Chat and Qwen1.5-72B-Chat exhibit similar performance, which may suggest a lack of training data in the psychology domain, preventing the models from leveraging their larger parameter counts. Many smaller Chinese models, such as Yi-6B-Chat and Qwen1.5-7B-Chat, outperform larger and stronger models like Mixtral-8x7B-Instruct, potentially due to the former having more training data in Chinese. Most models demonstrate lower performance in Psychological Statistics and Psychometrics compared to other subjects, implying these models need to enhance their reasoning abilities in psychology.
4.4 Chapter-level Performance
ConceptPsyalso provides chapter-level performance insights, offering more detailed information. From Figure 7, we can understand why the strong Qwen-72B-Chat model’s average score is lower than the smaller model Yi-34b-Chat. Qwen-72B-Chat underperforms Yi-34b-Chat in most chapters, particularly in reasoning chapters such as Applied Calculations in Statistical Psychology or Statistical Charts. In contrast, Mistral-8x7B-Instruct performs worse in theoretical chapters, possibly indicating a lack of corresponding knowledge in its training data. All models perform poorly in non-parametric tests, suggesting potential areas of improvement for developers. More chapter-level results can be found in Appendix E
5 Related Works
5.1 Chinese MMLU benchmarks
While the evolution of English language benchmarks continues to flourish (Hendrycks et al., 2020; Huang et al., 2023; Li et al., 2023; Zhong et al., 2023), Chinese MMLU benchmarks remain underdeveloped. The CLUE benchmark (Xu et al., 2020) is a widely used large-scale NLU benchmark for Chinese.The AGIEval benchmark (Zhong et al., 2023) expands this with questions from various Chinese exams, while MMCU (Zeng, 2023) includes questions from diverse domains. MCU (Zeng, 2023) expands on this by collecting questions from a broader range of domains. The C-EVAL benchmark Huang et al. (2023) gathers questions from different levels, including middle school, high school, and professional qualification exams, employing strategies to mitigate dataset leakage by using non-paper-based questions and simulation questions. ConceptMath (Wu et al., 2024) also evaluates LLM’s ability in a concept-wise manner. However, it differs from our work in that it (1) focuses on mathematics, (2) is not a comprehensive MMLU benchmark, and (3) is released more than three months after our work.
5.2 Benchmarks for LLMs in psychology
Although psychology is a crucial domain for achieving artificial intelligence, there are currently few benchmarks in this field. CMMLU has collected hundreds of questions as "professional psychology" to evaluate a model’s knowledge understanding and reasoning capabilities. Other works (Yang et al., 2023b; Amin et al., 2023; Lamichhane, 2023) approach this from a mental health perspective, utilizing various classification tasks, including various emotions, suicidal tendencies, and more. PsyEval (Jin et al., 2023) further expands in the mental health field by adding tasks such as Diagnosis Prediction.
6 Discussion
This paper introduces  ConceptPsy, the first comprehensive benchmark for manually collected psychology concepts in the psychology domain. This addresses the current issue where popular Chinese MMLU benchmarks either lack psychology subjects or have significant concept bias in psychology. In addition to providing an average score for each subject in  ConceptPsy, we manually label each question with a chapter-level concept to offer more fine-grained results. This assists developers in better improving and deploying their models. Furthermore, we conduct a case study on the presence of concept bias in current Chinese MMLUs, further underscoring the urgent need for  ConceptPsy.
Limitations
We have solely explored the concept-related bias within the Chinese MMLU series of datasets and have not investigated biases in other areas. Additionally, due to resource constraints, we have only collected and analyzed the concepts from limited subjects. Our work exclusively examines Chinese benchmarks and does not extend to benchmarks in other languages.
Ethics Statement
Appendix A Human Evaluation
@@
Appendix B Competition with Human-designed Questions
@@
Appendix C Chapter-level Statistic of ConceptPsy
@@
Appendix D Experiments Details of Calculating Concept Coverage Rate
@@
Appendix E More Results on Fine-grained Performance
@@
Appendix F Qualifications of Our psychology Annotators
Our psychological annotators is professional. We enlist three psychological annotators with diverse expertise. The first is a postdoctoral researcher specializing in experimental and counseling psychology. The second is a registered psychology researcher with the British Psychological Society with five years of research experience in interdisciplinary psychology laboratories. The third, a counseling psychologist, brings over three years of practical counseling experience.
@@ Appendix H Prompts for Questions Generation
Figure 13: multiple type prompt

### v4

@@ 4.3 Results
The main results of different LLMs on  ConceptPsy  are shown in Table 6. We also provide fine-grained results at Figure 7, 8 and 9.
Strong Chinese models achieve comparable overall results with GPT-4: In Table 6, surprisingly, some Chinese models achieve comparable performance close to that of GPT-4 on  ConceptPsy. Note that GPT-4 have the concepts as input during the question generation process but not in the evaluation stage. Overall, GPT-4 performed significantly better in subjects requiring more reasoning ability, such as Psychological Statistics (79% compared to 69%). However, in subjects requiring a deep understanding of the Chinese context and knowledge, such as Clinical & Counseling, Yi-34B-Chat performed better. This reflects the differences in abilities between the two models, indicating that GPT-4 might have insufficient training data in some areas of Chinese psychology, but its reasoning ability is much stronger.
Chinese models struggle with math and reasoning: Analysis from Figure 7, 8, and 9 reveals that Chinese models’ performance drops on concepts requiring math and reasoning, like “Mathematical Basis of Inferential Statistics” and “Non-Parametric Tests”. While these models excel in knowledge-based areas, they falter in math and reasoning segments. Mixtral-8x7B-Ins stands out by underperforming in knowledge questions but outshining others in math and reasoning. This suggests that despite high accuracy from training on extensive Chinese texts, Chinese models’ math and reasoning skills, especially in psychology, need enhancement.
Models show varied significant chapter performance: There’s notable variation in how models perform across chapters. Yi-34B-Chat, for example, despite leading in average accuracy, exhibits over 50% performance differences between chapters—excelling in basic concepts but lagging in math reasoning, even behind Mixtral-8x7B-Ins with its 10% lower average accuracy. This underscores  ConceptPsy’s role in offering detailed insights for model enhancement.
Variability in conceptual mastery across models from the same series: Although Qwen1.5-14B-Chat and Qwen1.5-72B-Chat achieve similar overall results, this does not mean their performance is similar across all chapters. As shown in Figures 7 and 8, Qwen1.5-72B-Chat performs much worse in chapters that focus on math and reasoning, like "Non-Parametric Tests" and "Parameter Estimation." While the overall accuracy suggests that both models perform similarly, the detailed results from  ConceptPsy  show that Qwen1.5-72B-Chat is generally better than Qwen1.5-14B-Chat in most chapters, except for a few math-related ones where it falls behind. This demonstrates how  ConceptPsy  can highlight specific strengths and weaknesses, helping developers make targeted improvements.
@@
4.4 Ablation Study
Human baseline. To assess  ConceptPsy’s ability to distinguish between human expertise levels, we ask professional counselors and non-professional grad students to evaluate a subset of tasks. Table 7 shows significant performance differences (Appendix A).
The effect of linguist style. To investigate the effect of different linguistic styles of generated questions, we paraphrase each question using Qwen-70b-Chat evaluate LLMs on it. The results are shown in Table 8. There is a slight performance drop on paraphrased questions, potentially due to the lack of manual review for these paraphrased versions. Nonetheless, the consistency in model rankings and overall performance suggests that for multiple-choice questions, stylistic variations of the questions have limited effect.
5 Related Works
Chinese MMLU Benchmarks: While English language benchmarks continue to evolve (Hendrycks et al., 2020; Huang et al., 2023; Li et al., 2023; Zhong et al., 2023), Chinese MMLU benchmarks remain underdeveloped. The CLUE benchmark (Xu et al., 2020) is a widely used large-scale NLU benchmark for Chinese. AGIEval (Zhong et al., 2023) expands this with questions from various Chinese exams, and MMCU (Zeng, 2023) includes questions from diverse domains. The C-EVAL benchmark Huang et al. (2023) gathers questions from different educational levels and professional exams, using non-paper-based and simulation questions to prevent dataset leakage. ConceptMath (Wu et al., 2024) evaluates LLMs in a concept-wise manner but focuses solely on mathematics, is not a comprehensive MMLU benchmark, and was released more than three months after our work.
Benchmarks for LLMs in Psychology. Although psychology is a crucial domain for achieving artificial intelligence, there are currently few benchmarks in this field. CMMLU has collected hundreds of questions as "professional psychology" to evaluate a model’s knowledge understanding and reasoning capabilities. Other works (Yang et al., 2023b; Amin et al., 2023; Lamichhane, 2023) approach this from a mental health perspective, utilizing various classification tasks, including various emotions, suicidal tendencies, and more. PsyEval (Jin et al., 2023) further expands in the mental health field by adding tasks such as “Diagnosis Prediction”, but lacking comprehensive subjects in college-level psychology.
6 Discussion
Domain benchmarks are crucial for adapting LLMs. However, a comprehensive benchmark for assessing models’ knowledge and reasoning abilities in psychology is lacking. To address this gap, we introduce  ConceptPsy. In designing  ConceptPsy, we reveal concept bias issues in previous Chinese MMLU benchmarks and study concept coverage rates and performance variances. To tackle these issues, we propose a concept-wise question generation framework and provide chapter-level fine-grained results. Evaluation across various models demonstrates  ConceptPsy’s effectiveness in highlighting model strengths and weaknesses.
Limitations
We focused exclusively on concept bias within the Chinese MMLU dataset series, not exploring biases in other dimensions due to the labor-intensive nature of collecting the required concepts in a discipline, which necessitates trained annotators. Our experiments were thus confined to the field of psychology. Although our question generation method could be easily applied to other domains, due to copyright issues and the absence of corresponding multi-choice questions for many concepts, we opted to use GPT-4 for question generation instead of relying on human-designed questions.
Ethics Statement
Appendix A Human Evaluation
@@
Appendix B The effect of Linguistic Style of generated questions
To investigate the effect of linguistic style, we paraphrase each question using Qwen-70b-Chat. We then evaluated the performance difference between its style and GPT-4’s style. The results are summarized in the table below:
Clinical &
Counseling
Psy of
Personality
Abnormal
Psy
History of
Psy
General
Psy
Psy-
chometrics
Social
Psy
Management
Psy
Psychological
Statistics
Experimental
Psy
Developmental
Psy
Educational
Psy
Avg
Mistral-7B-Instruct 0.71 0.63 0.64 0.58 0.57 0.59 0.69 0.71 0.53 0.64 0.62 0.69 0.63
gpt-3.5-Turbo 0.80 0.72 0.79 0.67 0.70 0.68 0.77 0.78 0.60 0.73 0.71 0.76 0.73
Yi-6B-Chat 0.80 0.75 0.87 0.67 0.80 0.70 0.83 0.83 0.58 0.69 0.76 0.86 0.76
Yi-34B-Chat 0.86 0.80 0.90 0.78 0.83 0.78 0.87 0.88 0.70 0.82 0.82 0.85 0.82
Table 8: Performances of models on questions paraphrased by Qwen-70b-Chat
In comparison to Table 6, models perform slightly worse on paraphrased questions generated by Qwen-70b-Chat. This may be because the paraphrased questions are not manually reviewed. However, despite the style changes, the model rankings and overall performance remained consistent. This indicates that for multiple-choice questions, the style of the questions has minimal impact on evaluation.
@@ Appendix B The effect of Linguistic Style of generated questions
Appendix C Competition with Human-designed Questions
@@
Appendix D Chapter-level Statistic of ConceptPsy
@@
Appendix E Experiments Details of Calculating Concept Coverage Rate
@@
Appendix F More Results on Fine-grained Performance
@@ Appendix H Prompts for Questions Generation
Figure 13: Multiple types prompt